the mean monthly Internet bill is $32.79 per household. A sample of 50 households in the Belize District showed a sample mean of $30.63. Use a population standard deviation $5.60.

  

 What is the value of the test statistic? 

c) What is the p-value?

d) At α = .05, what is your conclusion?


1
Expert's answer
2022-04-28T08:41:26-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=32.79"

"H_1:\mu\not=32.79"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

b) The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}"

c) The p-value is "p=2P(Z<z), z\le0" or "p=2P(Z>z), z\ge0."


d) Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."


"z=\dfrac{30.63-32.79}{5.60/\sqrt{50}}\approx-2.7274"

Since it is observed that "|z| = 2.7274 >1.96= z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-2.7274)=0.006384," and since "p=0.006481<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 32.79, at the "\alpha = 0.05" significance level.